Obtain the equation of a parabola with focus (4,-6) and directrix as y=-2
1
Expert's answer
2021-09-13T00:26:20-0400

A parabola is defined as follows:

For a given point, called the focus, and a given line not through the focus, called the directrix, a parabola is the locus of points such that the distance to the focus equals the distance to the directrix.

Let "P(x,y)" be the point on the parabola. Then


"(x-4)^2+(y+6)^2=(y-(-2)^2"

"(x-4)^2=-y^2-12y-36+y^2+4y+4"

"(x-4)^2=-8y-32"

Therefore the required equation of the parabola is


"(x-4)^2=-8y-32"